A car leaves Beirut at 12h and reaches said a at 12h 37min the distance covered is 45000m. 1)calculate duration of trip in SI unit. 2) calculate average speed of the car 3) The car continues its motion to south of Lebanon and attains a village(k) at 13h16min . Calculate the distance covered between saida and village(k) knowing that the avarage speed of the car between saida and village is equal to that calculated in part 2.
1
Expert's answer
2020-03-17T10:02:05-0400

1)


"t=37(60)=2220\ s"

2)


"v=\frac{d}{t}=\frac{45000}{2220}=20.27\frac{m}{s}"

3)


"t'=(60+16-37)60=2340\ s"

"d'=vt'=(20.27)(2340)=47430\ m"